# Autograding Mathematical Induction Proofs with Natural Language Processing

### v1

@@ Abstract
In mathematical proof education, there remains a need for interventions that help students learn to write mathematical proofs. Research has shown that timely feedback can be very helpful to students learning new skills. While for many years natural language processing models have struggled to perform well on tasks related to mathematical texts, recent developments in natural language processing have created the opportunity to complete the task of giving students instant feedback on their mathematical proofs. In this paper, we present a set of training methods and models capable of autograding freeform mathematical proofs by leveraging existing large language models and other machine learning techniques. The models are trained using proof data collected from four different proof by induction problems. We use four different robust large language models to compare their performances, and all achieve satisfactory performances to various degrees. Additionally, we recruit human graders to grade the same proofs as the training data, and find that the best grading model is also more accurate than most human graders. With the development of these grading models, we create and deploy an autograder for proof by induction problems and perform a user study with students. Results from the study shows that students are able to make significant improvements to their proofs using the feedback from the autograder, but students still do not trust the AI autograders as much as they trust human graders. Future work can improve on the autograder feedback and figure out ways to help students trust AI autograders.
@@ 1 Introduction
Writing mathematical proofs has been identified as an important [1, 2, 3] and yet challenging topic [4] in computing education and mathematics education. A large body of research has shown that timely feedback is crucial to student learning [5, 6]. However, students are largely unable to receive timely feedback on written proofs due to the need to have proofs collected and hand-graded by instructors or teaching assistants. The ability to grade student proofs fully automatically with natural language processing (NLP) alleviates this need by allowing us to give students instant feedback on their proofs to let students iteratively enhance the quality of their proofs.
@@ 2.1 Research on Mathematical Proof Education
While there has been extensive work both on software tooling and its evaluation for helping students learn school geometry proofs [5], there is still a need for tools to provide students with interventions and scaffolding on learning to write college-level mathematical proofs. Stylianides and Stylianides’ review of the literature on teaching and learning proofs concluded that “more intervention-oriented studies in the area of proof are sorely needed”[8]. Several tools have been created which provide a visual method for students to construct simple mathematical proofs [9, 10, 11]. While these tools can act as scaffolding for students in the early stages of learning to write proofs, there are currently no existing software tools that can provide help to students working on the authentic task of writing proofs on their own from scratch. The grading models contributed by this paper enables to the creation of software tools which can give students automated feedback on their work as they practice writing natural language mathematical proofs on their own.
@@ 2.2 Natural Language Processing (NLP)
As a subfield of artificial intelligence, NLP primarily focuses on training and using models focusing on interpreting and generating natural languages used by humans. In 2017, NLP researchers invented the Transformer architecture, utilizing attention as its only mechanism and achieving good performance in learning the dependencies in inputs and outputs [12]. It also enabled easier training and inference by lowering the need of advanced training equipments or extended training time [12]. Further developments have led to the introduction of various pretrained large language models accessible such as BERT, Llama and Llama2, and OpenAI’s GPT family of models  [13, 14, 15]. In addition to being able to generate text, these models can take texts as inputs and return vectors representing the words or sentences in the texts. These vectors are called embeddings.
@@ 2.3 Automatic Short Answer Grading (ASAG)
Researchers have attempted to use sentence embeddings to perform ASAG. Various NLP models have been tested on their performance, with attention-based models performing relatively well in terms of accuracy, latency, reproducibility, and ease in fine-tuning [18]. Existing training data sets cover topics in data structures, introductory statistics, and a few other topics, with no coverage of mathematical proof problems [18]. Previous work has also produced systems able to grade student’s descriptions of short snippets of Python code, achieving over 85% accuracy with a dataset of less then 600 samples [19]. These models have successfully been deployed in the classroom to help students receive timely feedback and reduce instructor grading load [20].
@@
The focus on ASAG has been put into mathematical contexts as well, with limited success so far. Researchers have developed Mathematical Language Processing to perform mathematical ASAG, and achieved an absolute mean error of $0.04$ out of a full grade of $3$ [21]. However, the approach taken by Mathematical Language Processing only allows a small set of mathematical vocabulary, and relies heavily on symbolic manipulation on the answers. As a result, it will not be able to autograde mathematical proofs, which usually involves reasoning using natural language, such as proof by induction or proof on graphs. MathBERT was created by fine-tuning BERT, aimed at dealing with a larger mathematical vocabulary and more complex mathematical contexts [22]. It achieved over $90\%$ accuracy on autograding tasks, but the problems used were one-sentence Q&A problems that did not require much reasoning or formula manipulation, and were much shorter than the full inductive proofs that we treat in this work. To deal with verifying longer proofs, researchers have also attempted autoformalization [23], translating natural language into formal logic. This should make checking the correctness of the steps in a proof easier. However, even with several years of research, none of the existing work is able to translate even half of the proofs to formal logic [23]. The low success rate makes autoformalization not suited for grading mathematical proofs at the moment.
@@
To our knowledge, no tools have successfully dealt with grading mathematical proofs using natural language processing. More emphasis needs to be put on verifying the natural language reasoning part of the proofs in addition to the arithmatic formulae.
@@
Utilizing ASAG in practice poses a few difficulties. First of all, it is hard for the ASAG systems to be perfect: mistakes will likely happens during the grading process. In the context of potentially inaccurate feedback from the graders, helping students navigate through the feedback can be crucial in achieving the full potentials from ASAG systems [24]. Another challenge facing the wider use of ASAG is students’ trust in AI. Previous research around AI trust has shown that people are more likely to trust human decisions than algorithmic decisions, especially when the tasks are more subjective [25, 26]. Some previous work has observed that students underestimated the accuracy of some ASAG systems, believing they were graded incorrectly when the grades were accurate [27]. The lack of trust in the technology behind the ASAG systems could hinder the effectiveness of the ASAG systems.
@@ 3.1 Data
The proofs used in this study were collected from the pretest and posttest of a series of studies designed to measure the learning gains of students using Proof Blocks [28, 29, 30]. The study participants were students recruited from a Discrete Mathematics course at the University of Illinois Urbana-Champaign. The proofs were originally collected in the form of markdown files, and later graded by members of the research team. The proofs were graded in a multi-step process which involved multiple members of the research team grading the proofs independently, and then meeting together to discuss disagreements. The graders were able to reach an inter-rater reliability rating of Cronbach’s $\alpha$ at 0.82, 0.88, and 0.92 for each of the three studies respectively [28, 29, 30].
These measurements show very high consistency among the agreement between researchers, and so we use these labels as ground truth for our training and further analysis. Since the grading of informally written mathematical proofs can be subjective, we recognize that our AI autograders and the human baseline comparison must not only agree with fundamental truths but also align with the grading standards set by the research team that created the original labels. However, we note that this is the same task that an ASAG system or a teaching assistant must accomplish when grading student work for a course—not only verifying underlying truth but also adhering to the grading criteria set by the instructor, who acts as the ground truth in this context. We also believe that our use of only technical, and not stylistic rubric points, as well as the rigorous process we used to have multiple researchers agree on a grading scheme (rather than have them decided by a single instructor), make our grading system as objective as possible in this context.
The proofs were graded using seven rubric points, each identifying a step of the induction proof:
@@
Figure 1: An example proof for P2. In the original labeling by previous researchers, this proof was labeled correct on rubric points R1: Identifying the base case, R2: Proving the base case, and R6: Breaking down the inductive step, and incorrect on all other rubric points.
@@ 3.2 Model selection for Embeddings
Several pretrained large language models are capable of calculating embeddings such as BERT and its mathematically fine-tuned version MathBERT [22], Llama2 and its mathematically fine-tuned version Llemma [31], and GPT-3 [32]. Base BERT and base Llama2 are excluded because MathBERT and Llemma have been observed to perform better than their corresponding base models when dealing with mathematical contexts [22, 31]. MathBERT is among the best pretrained language models at mathematical language, and runs on a CPU. The memory requirement is typically less than 600MB. Because of the low memory requirement, MathBERT allows for cheap deployment. It is worthwhile trying MathBERT despite having fewer parameters than most large language models, because if MathBERT can achieve similar performances as the other large language models, it would also be cheaper to deploy. This can be a determining factor of the grading models if they need to be scaled in the future to serve more students. GPT-3 is the best accessible language model at the moment for embeddings. We use the latest embedding endpoint text-embedding-3-large for analysis, as it outperforms the old endpoint text-embedding-ada-002 by 0.3% in accuracy in our grading tasks. GPT-3 requires API calls for embeddings, and each API call costs a small amount of money. This additional cost should be taken into consideration when scaling this grading model for larger number of students. We did not use GPT-4 as it does not yet support embeddings, which are preferred to text completion for classification tasks such as ours [32]. Llemma has two versions, Llemma7b with 7 billion parameters, and Llemma34b with 34 billion parameters. Both can be run locally, but take a significant amount of RAM and require a GPU even for inference. In our study, we utilize MathBERT, GPT-3, Llemma34b, and Llemma7b for embeddings and then compare the performances of the models.
@@ 3.3 Model fitting
The dataset is broken down into training, testing, and validation sets, with $15\%$ for testing and another $15\%$ for validation. The remaining $70\%$ are used for training. Performance on the test set will be used to compare the performances of models. The $15\%$ data for validation is reserved for use in future work when the model complexity is increased.
In training our grading model, we chose a batch size of $128$ as it is a common choice for training networks. Number of epochs for training range from $100$ to $1000$, with a step size of $100$ epochs. Initially, the learning rate is set to 00 and linearly increased to a peak value of $0.001$ over the first $60\%$ of the training epochs. This linear increase provides a warm-up period for the model to gradually adapt to the task. Subsequently, during the second half of the training, the learning rate is exponentially decreased to $\frac{1}{10}$ of the peak value, allowing the model to fine-tune its parameters and converge towards a stable solution.
To save computational resources, we generate all embeddings by running the pretrained models once on the entire proof dataset. Then we save the mapping between each proof and its corresponding embedding to a pickle file. The two Llemma models were run on our University’s supercomputing cluster, using eight NVIDIA A100 GPUs simultaneously, taking about one second on average to turn one proof into embeddings; GPT-3 embeddings were retrieved from an OpenAI API call; MathBERT was run on a single NVIDIA RTX4090 GPU. The customized grading models are built and trained using PyTorch [33] on the same NVIDIA GPU. For each model, trainings completed at a rate of about 2 seconds per 100 training epochs and 10 batchs.
@@ 3.4 Comparison to human graders
Theoretically, it is hard for machine learning models to perfectly grade the freeform mathematical proofs. However, it can also be hard for humans to agree on the grading of mathematical proofs both due to ambiguity in language and honest mistakes. In the case that there is any subjectivity in the grading, human graders also need to match the specification set by the course instructor. To our knowledge, no prior study exists that quantifies the accuracy of humans grading mathematical proofs to match an instructor specification. We are also interested in how well human graders can match the grading specification compared with our NLP grading models.
Using the last author’s professional network, we recruited graduate students from several universities who had been teaching assistants in discrete mathematics courses. These graduate students all had experience grading mathematical induction proofs. 9 graduate students from 4 different institutions participated in the grading tasks. For each human grader, we sent 15 student proofs from each of the 4 proof problems (for a total of 60 problems each), and asked them to grade these proofs using the same rubric. Along with the proofs to grade, we also provided 15 examples of graded proofs for each of the problems, the original problem statements, and detailed notes with explanations of the rubrics created by previous researchers. This combination of example labeled proofs and detailed rubric explanations were designed to set the graders up to be as successful as possible on their grading task. In fact, we provided the graders with more training materials for grading than many instructors do for grading for actual courses (in our experience). The graders were also free to ask us any questions regarding problems encountered during the grading process, as they would be when grading student work for an actual course. Each grader received a \$100 gift card as compensation for their grading efforts. This human subjects data collection procedure was approved by the institutional review board at Utah State University.

### v2

@@ Abstract
In mathematical proof education, there remains a need for interventions that help students learn to write mathematical proofs. Research has shown that timely feedback can be very helpful to students learning new skills. While for many years natural language processing models have struggled to perform well on tasks related to mathematical texts, recent developments in natural language processing have created the opportunity to complete the task of giving students instant feedback on their mathematical proofs. In this paper, we present a set of training methods and models capable of autograding freeform mathematical proofs by leveraging existing large language models and other machine learning techniques. The models are trained using proof data collected from four different proof by induction problems. We use four different robust large language models to compare their performances, and all achieve satisfactory performances to various degrees. Additionally, we recruit human graders to grade the same proofs as the training data, and find that the best grading model is also more accurate than most human graders in terms of grading accuracy. With the development of these grading models, we create and deploy an autograder for proof by induction problems and perform a user study with students. Results from the study shows that students are able to make significant improvements to their proofs using the feedback from the autograder, but students still do not trust the AI autograders as much as they trust human graders. Future work can improve on the autograder feedback and figure out ways to help students trust AI autograders.
@@ 1 Introduction
Writing mathematical proofs has been identified as an important [1, 2, 3] and yet challenging topic [4] in computing and mathematics education. A large body of research has shown that timely feedback is crucial to student learning [5, 6]. However, students are largely unable to receive timely feedback on written proofs due to the need to have proofs collected and hand-graded by instructors or teaching assistants. The ability to grade student proofs fully automatically with natural language processing (NLP) alleviates this need by allowing us to give students instant feedback on their proofs to let students iteratively enhance the quality of their proofs.
@@ 2.1 Research on Mathematical Proof Education
It has been acknowledged that teaching and using proofs plays an important role in mathematics education [8]. While there has been extensive work both on software tooling and its evaluation for helping students learn school geometry proofs [5], there is still a need for tools to provide students with interventions and scaffolding on learning to write college-level mathematical proofs. Stylianides and Stylianides’ review of the literature on teaching and learning proofs concluded that “more intervention-oriented studies in the area of proof are sorely needed”[9]. There have been interventions that focus on helping students understand the need for proofs, as students often believe empirical arguments without seeing the need for proof [9, 10, 11, 12]. There are also various experience reports on novel interventions that instructors have tried with limited evidence on their effectiveness [13, 14, 15]. Several tools have been created which provide a visual method for students to construct simple mathematical proofs [16, 17, 18]. While these tools can act as scaffolding for students in the early stages of learning to write proofs, there are currently no existing software tools that can provide help to students working on the authentic task of writing proofs on their own from scratch.
Within the field of mathematical proofs, proof by induction is a particular point of interest because of its requirement for various cognitive schemas and abstraction skills [19, 20]. Its mostly non-explanatory nature [21] often leaves students uncertain about when to chose induction over other explanatory proof techniques. Previous research has explored methods for teaching students both when to use induction [22], and how to construct induction proofs correctly [20, 23]. Several misconceptions have been identified during the process [24, 23, 15, 25]. In Section 3.1, we will discuss how the items in our rubric align with many of the misconceptions reported in the literature. A more exhaustive list of all misconceptions can be found in recent work in mathematics education [15, 26]. Most of the misconceptions listed in prior work have also shown up in our own training data set [26]. Despite the prevalence of these issues, students currently lack tools to help them address these misconceptions and practice constructing induction proofs effectively. The grading models introduced in this paper enable the development of software tools that provide students with automated feedback on their work as they practice writing natural language mathematical proofs on their own.
@@ 2.2 Natural Language Processing (NLP)
As a subfield of artificial intelligence, NLP primarily focuses on training and using models to interpret and generate natural languages. In 2017, NLP researchers developed the Transformer architecture, which relies solely on attention mechanisms, achieving good performance in learning the dependencies in inputs and outputs [27]. It also enabled easier training and inference by lowering the need of advanced training equipments or extended training time [27]. Further developments have led to the introduction of various pretrained large language models accessible such as BERT, Llama and Llama2, and OpenAI’s GPT family of models  [28, 29, 30]. In addition to being able to generate text, these models can take texts as inputs and return vectors representing the words or sentences in the texts. These vectors are called embeddings.
@@ 2.3 Automatic Short Answer Grading (ASAG)
Researchers have attempted to use sentence embeddings to perform ASAG. Various NLP models have been tested on their performance, with attention-based models performing relatively well in terms of accuracy, latency, reproducibility, and ease in fine-tuning [33]. Existing training data sets cover topics in data structures, introductory statistics, and a few other topics, with no coverage of mathematical proof problems [33]. One thread of previous work utilized BERT in automatic scoring of reading comprehension by fine-tuning BERT with a set of reading comprehension problems and answers, achieving accuracy almost as good as human graders [34]. However, it lacks the ability to easily transfer to new problems, creating the need to fine-tune the model repeatedly when new problems are added, which can be a computationally expensive process. The same drawback would apply to other ASAG attempts utilizing fine-tuning as the primary way of grading. Previous work has also produced systems able to grade student’s descriptions of short snippets of Python code by creating item-specific models, achieving over 85% accuracy with a dataset of less then 600 samples [35]. These models have successfully been deployed in the classroom to help students receive timely feedback and reduce instructor grading load [36].
@@
Efforts to apply ASAG in mathematical contexts have been made but have achieved only limited success to date. Researchers have developed Mathematical Language Processing to perform mathematical ASAG, and achieved an absolute mean error of $0.04$ out of a full grade of $3$ [37]. However, the approach taken by Mathematical Language Processing allows only a limited set of mathematical vocabulary, and relies heavily on symbolic manipulation on the answers. As a result, it cannot autograde mathematical proofs, which usually involves reasoning using natural language, such as proof by induction or proof on graphs. MathBERT was created by fine-tuning BERT, aimed at dealing with a larger mathematical vocabulary and more complex mathematical contexts [38]. It achieved over $90\%$ accuracy on autograding tasks. However, the problems addressed were one-sentence Q&A tasks that required minimum reasoning or formula manipulation. These tasks were significantly shorter than the full inductive proofs considered in this work. To deal with verifying longer proofs, researchers have also attempted autoformalization [39], translating natural language into formal logic. This should make checking the correctness of the steps in a proof easier. However, even with several years of research, none of the existing work is able to translate even half of the proofs to formal logic [39]. The low success rate makes autoformalization not suited for grading mathematical proofs at the moment.
@@
To our knowledge, no tools have successfully dealt with grading mathematical proofs using natural language processing. More emphasis needs to be put on verifying the natural language reasoning part of the proofs in addition to the arithmetic formulae.
@@
Utilizing ASAG in practice poses a few difficulties. First of all, it is hard for the ASAG systems to be perfect: mistakes will likely happen during the grading process. In the context of potentially inaccurate feedback from the graders, helping students navigate through the feedback can be crucial in achieving the full potential from ASAG systems [40]. Another challenge facing the wider use of ASAG is students’ trust in AI. Previous research around AI trust has shown that people are more likely to trust human decisions than algorithmic decisions, especially when the tasks are more subjective [41, 42]. Some previous work has observed that students underestimated the accuracy of some ASAG systems, believing they were graded incorrectly when the grades were accurate [43]. The lack of trust in the technology behind ASAG systems could hinder their effectiveness.
@@ 3.1 Data
The proofs used in this study were collected from the pretest and posttest of a series of studies designed to measure the learning gains of students using Proof Blocks [44, 45, 46]. The participants were students recruited from a Discrete Mathematics course at the University of Illinois Urbana-Champaign. The proofs were originally collected in the form of markdown files, and later graded by members of the research team. The proofs were graded in a multi-step process which involved multiple members of the research team grading the proofs independently, and then meeting together to discuss disagreements. The graders were able to reach an inter-rater reliability rating of Cronbach’s $\alpha$ at 0.82, 0.88, and 0.92 for each of the three studies respectively [44, 45, 46].
These measurements show very high consistency in agreement between researchers, allowing us to use these labels as ground truth for training and further analysis. Since the grading of informally written mathematical proofs can be subjective, we recognize that our AI autograders and the human baseline comparison must not only agree with fundamental truths but also align with the grading standards set by the research team that created the original labels. However, we note that this is the same task that an ASAG system or a teaching assistant must accomplish when grading student work for a course—not only verifying underlying truth but also adhering to the grading criteria set by the instructor, who acts as the ground truth in this context. We also believe that our use of only technical, and not stylistic rubric points, as well as the rigorous process we used to have multiple researchers agree on a grading scheme (rather than have them decided by a single instructor), make our grading system as objective as possible in this context.
The proofs were graded using seven rubric points, each identifying a step of the induction proof.
@@
Each of these rubric points was designed to address specific aspects of constructing a complete and correct induction proof. The rubric points also align with common misconceptions identified in prior research. Students’ difficulties in identifying and proving base cases (R1 and R2) were discussed by Norton et al. [15], who also reference earlier literature documenting these issues. Many students have difficulties in stating the inductive assumption correctly. For example, some students will assume the entire conjecture to be true rather than the inductive implication [15, 19]. Students will also fail to give the correct upper or lower bound for the inductive implication [26]. These issues are captured by rubric points R3 and R4. The rubric points R5, R6, and R7 correspond to the misconception “Sets up induction proof properly but does not succeed in proving that $P(n)\Rightarrow P(n+1)$" documented by Dubinsky [23]. The grading instructions included in our public dataset [7] give detailed instructions on when each rubric point should be marked as correct for a given proof, including examples.
@@
Figure 1: An example proof for P2. In the original labeling by previous researchers, this proof was labeled correct on rubric points R1 (Identifying the base case), R2 (Proving the base case), and R6 (Breaking down the inductive step), and incorrect for all other rubric points.
@@ 3.2 Model selection for Embeddings
Several pretrained large language models are capable of calculating embeddings such as BERT and its mathematically fine-tuned version MathBERT [38], Llama2 and its mathematically fine-tuned version Llemma [47], and GPT-3 [48]. Base BERT and base Llama2 are excluded because MathBERT and Llemma have been observed to perform better than their corresponding base models when dealing with mathematical contexts [38, 47]. MathBERT is among the best pretrained language models at mathematical language, and runs on a CPU. The memory requirement is typically less than 600MB. Because of the low memory requirement, MathBERT allows for cheap deployment. It is worthwhile trying MathBERT despite having fewer parameters than most large language models, because if MathBERT can achieve similar performances as the other large language models, it would be cheaper to deploy. This could become a critical factor when selecting the grading models if they need to be scaled in the future to accommodate a larger number of students. GPT-3 is the best accessible language model at the moment for embeddings. We use the latest embedding endpoint text-embedding-3-large for analysis, as it outperforms the old endpoint text-embedding-ada-002 by 0.3% in accuracy in our grading tasks. GPT-3 requires API calls for embeddings, and each API call costs a small amount of money. This additional cost should be taken into consideration when scaling this grading model for larger number of students. We did not use GPT-4 as it does not yet support embeddings, which are preferred to text completion for classification tasks such as ours [48]. Llemma has two versions, Llemma7b with 7 billion parameters, and Llemma34b with 34 billion parameters. Both can be run locally, but take a significant amount of RAM and require a GPU even for inference. In our study, we utilize MathBERT, GPT-3, Llemma34b, and Llemma7b for embeddings and then compare the performances of the models.
@@ 3.3 Model fitting
The dataset is broken down into training, testing, and validation sets, with $15\%$ for testing and another $15\%$ for validation. The remaining $70\%$ are used for training. Performance on the test set will be used to compare the performances of models. The $15\%$ data for validation is reserved for future work with increased model complexity.
For the training of our grading model, we chose a batch size of $128$ as it is a common choice for training networks. Number of epochs for training range from $100$ to $1000$, with a step size of $100$ epochs. Initially, the learning rate is set to 00 and linearly increased to a peak value of $0.001$ over the first $60\%$ of the training epochs. This linear increase provides a warm-up period for the model to gradually adapt to the task. Subsequently, during the second half of the training, the learning rate is exponentially decreased to $\frac{1}{10}$ of the peak value, allowing the model to fine-tune its parameters and converge towards a stable solution.
To save computational resources, we generate all embeddings by running the pretrained models once on the entire proof dataset. Then we save the mapping between each proof and its corresponding embedding to a pickle file. We ran the two Llemma models on our University’s supercomputing cluster, using eight NVIDIA A100 GPUs simultaneously, taking about one second on average to turn one proof into embeddings; GPT-3 embeddings were retrieved from an OpenAI API call, and we ran MathBERT on a single NVIDIA RTX4090 GPU. The customized grading models were built and trained using PyTorch [49] on the same NVIDIA GPU. For each model, trainings completed at a rate of about 2 seconds per 100 training epochs and 10 batchs.
@@ 3.4 Comparison to human graders
Theoretically, it is hard for machine learning models to perfectly grade the freeform mathematical proofs. However, it can also be hard for humans to agree on the grading of mathematical proofs both due to ambiguity in language and honest mistakes. In the case that there is any subjectivity in the grading, human graders also need to match the specification set by the course instructor. To our knowledge, no prior study has quantified the accuracy of humans grading mathematical proofs to match an instructor specification. Additionoally, we want to compare how closely human graders can match the grading specification relative to our NLP grading models.
Using the last author’s professional network, we recruited graduate students from several universities who had been teaching assistants in discrete mathematics courses. These graduate students all had experience grading mathematical induction proofs. A total of 9 graduate students from 4 different institutions participated in the grading tasks. For each human grader, we sent 15 student proofs from each of the 4 proof problems (for a total of 60 problems each), and asked them to grade these proofs using the same rubric. Along with the proofs to grade, we also provided 15 examples of graded proofs for each of the problems, the original problem statements, and detailed notes with explanations of the rubrics created by previous researchers. This combination of example labeled proofs and detailed rubric explanations were designed to set the graders up to be as successful as possible on their grading task. In fact, we provided the graders with more training materials for grading than many instructors do for grading of actual courses (in our experience). The graders were also free to ask us any questions regarding problems encountered during the grading process just as they would be when grading student work in a real course setting. Each grader received a \$100 gift card as compensation for their grading efforts. This human subjects data collection procedure was approved by the institutional review board at Utah State University.